The length of a rectangle is 1m less than twice its width, and the area of the rectangle is 36m^2. Find the dimensions of the rectangle.






Answer by jim_thompson5910(35256)

x(2x - 1) = 36

2x^2 - x = 36

2x^2 - x - 36 = 0

(2x - 9)(x + 4) = 0

2x-9 = 0 or x+4 = 0

2x = 9 or x = -4

x = 9/2 or x = -4

x = 4.5 or x = -4

Toss out the negative solution.

So x = 4.5, 2x-1 = 2(4.5)-1 = 8

Therefore, the rectangle has a length of 8 m and a width of 4.5 m